OpenStudy (alimameh):

Suppose f:[−1,1]→R is a function with the property that x2+f(x)2=1 and suppose f(3/5)=−4/5. Use implicit differentiation to compute f′(3/5).

OpenStudy (3mar):

What did you get? Any ideas?

OpenStudy (alimameh):

Since x2+f(x)2=1 holds on an open interval around x=3/5, we can differentiate both sides to discover something about f′(3/5). ddx(x2+f(x)2)=2x+2f(x)f′(x). On the other hand, ddx(1)=0. Consequently, 2x+2f(x)f′(x)=0 for values of x around 3/5. Plugging in x=3/5 yields 2⋅34+2f(34)f′(34)=0. Solving, we find that f′(34)=−(3/5)/f(3/5). But f(3/5)=−4/5. Therefore f′(34)=−(3/5)/(−4/5)=3/4.

OpenStudy (holsteremission):

The idea at least is correct, but there seem to be typos throughout. \[\frac{\mathrm d}{\mathrm dx}\left(x^2+f(x)^2\right)=\frac{\mathrm d}{\mathrm dx}1\implies 2x+f(x)f'(x)=0\]and so plugging \(x=\dfrac{3}{5}\) and using the fact that \(f\left(\dfrac{3}{5}\right)=-\dfrac{4}{5}\), you get \[2\left(\frac{3}{5}\right)+\left(-\frac{4}{5}\right)f'\left(\frac{3}{5}\right)=0\implies f'\left(\frac{3}{5}\right)=\frac{\frac{6}{5}}{\frac{4}{5}}=\frac{3}{2}\]